OpenStudy (kidjov):

The expression -3m - [2m + (5 - m)] + 7 was simplified as 2 - 4m. Without simplifying, explain how you can show that it has been simplified correctly.

OpenStudy (kidjov):

i really need this question answered :( i dont know it

OpenStudy (skullpatrol):

Can you simplify the expression?

OpenStudy (kidjov):

its already simplify into 2 - 4m

OpenStudy (skullpatrol):

How do you know that is right?

OpenStudy (kidjov):

its part of the question.

OpenStudy (skullpatrol):

That's true. But did you check?

OpenStudy (welshfella):

You'll know if its right if you go through the process of simplification correctly

OpenStudy (welshfella):

for example did you distribite the negative over the paretnheses correctly? Did all the signs change in the parentheses

OpenStudy (welshfella):

did dropping the inside parentheses make any difference to the sign of 5 and -m ?

OpenStudy (welshfella):

- did you do that part first?

OpenStudy (skullpatrol):

One way to check is to let m=0 in the expression: -3m - [2m + (5 - m)] + 7 and see if you get the same number if you let m=0 in 2 - 4m.

OpenStudy (skullpatrol):

By testing different values in each you can show they are equivalent ie the simplification was done correctly.

OpenStudy (skullpatrol):

The expressions -3m - [2m + (5 - m)] + 7 and 2 - 4m are equivalent if they have the same solution set for all values of the variable

OpenStudy (skullpatrol):

Let y= -3m-[2m + (5 - m)] + 7 and z= 2 - 4m. Now substitute 0 for m into both and see if y =?= z. If they are equivalent equations it will work for any value of m.

OpenStudy (skullpatrol):

-3m - [2m + (5 - m)] + 7 -3*0 - [2*0 + (5 - 0)] + 7 = ? and 2 - 4m 2 - 4*0 = ??

OpenStudy (skullpatrol):

But the only sure way of knowing that the simplification has been done correctly is to set the two expressions equal to each other and then see if it reduces to 0=0. $$\Huge -3m - [2m + (5 - m)] + 7 = 2 - 4m$$

OpenStudy (kidjov):

thanks!

OpenStudy (skullpatrol):

Thanks for asking :-)